1. I need help with the following questions. Determine whether the point(1,4) is a solution of the system. Using the subsituting method.
6x + 9y = 42, 5x + 0y = 5
2. The point 1, 9 lines on the line 2x +10y =92. Determine if this a solution.

2x +10y = 92, -6x -2y = -24
3. Use the elimination method. If the system has no solution or an infinite number of solution state this.
-6x -6y = -12, -6x - 7y = 14
4. Use elimination method. If the system has no solution or infinite number of solution state this.
22x + 10y =147, 0x + 2y = -14

Answer by Alan3354(69443)

1. I need help with the following questions. Determine whether the point(1,4) is a solution of the system. Using the subsituting method.
6x + 9y = 42, 5x + 0y = 5
What sort of help? Sub x and y into the 2 eqns, and see if they're true.
----------------
2. The point (1,9) lines on the line 2x +10y =92. Determine if this a solution.
2*1 + 10*9 =? 92 Does it?
-----------------
-6x -2y = -24
-6*1 - 2*9 =? -24 Does it?
--------------------------
3. Use the elimination method. If the system has no solution or an infinite number of solution state this.
-6x - 6y = -12
-6x - 7y = 14
------------- Subtract
y = -26 --> 1 solution
-----------------------
4. Use elimination method. If the system has no solution or infinite number of solution state this.
22x + 10y =147
0x + 2y = -14 x5 --> 0x + 10y = -70
22x + 10y =147
0x + 10y = -70
----------------- Subtract
0 = 217 --> no solution